OpenStudy (kimberly_pr):

At what point do the lines y = 8x + 1 and 4x + 6y = 6 intersect?

OpenStudy (wolf1728):

y = 8x + 1 and 4x + 6y = 6 First, change the equations to: a) 8x -y = -1 b) 4x +6y = 6

OpenStudy (wolf1728):

Then multiply equation "a" by 6 a) 48x -6y = -6

OpenStudy (wolf1728):

Now add the 2 equations together a) 48x -6y = -6 b) 4x +6y = 6

OpenStudy (wolf1728):

What do you get?

OpenStudy (kimberly_pr):

Thank you :) can you help me with another question? I have to hurry

OpenStudy (kimberly_pr):

What is the product of the polynomials -x2y + 3xy – 8y and 4xy – 3y?

OpenStudy (wolf1728):

Okay go ahead

OpenStudy (kimberly_pr):

^^ the polynomials question

OpenStudy (wolf1728):

I'd do the multiplications one by one - (6 in all)

OpenStudy (wolf1728):

(-x2y + 3xy – 8y) * (4xy – 3y) = -x2y*4xy + (3xy*4xy) -(8y*4xy) + x^2*3y -3xy*-3y +8y*3y

OpenStudy (wolf1728):

Can you complete those?